### v1

@@ 1 Introduction
In the equilibrium setting, progress has been made in related inhomogeneous models, such as the gradient XX chain studied in Refs. [6, 7, 8], which allows for an exact characterization of local observables and entanglement. In contrast, when the linear potential acquires a time-dependent slope, the physics becomes more involved. Some special cases are known—such as quenches in the gradient XX chain [8], traveling impurity [9], or periodic driving [10, 11, 12, 13, 14, 15]. However, while hydrodynamic and numerical approaches have provided valuable insights [16, 17], a general exact solution for the dynamics of such driven Stark-localized systems has not been fully established. This limitation is especially relevant given the growing interest in quantum control [18, 19, 20], Floquet engineering [21, 22], and transport [23], where understanding the interplay between the driving and the resulting nonequilibrium dynamics is crucial.
More generally, and even restricting to non-interacting systems, exact results in the presence of driving are remarkably scarce. Notable exceptions include the case of a hardcore Bose gas in a time-dependent harmonic trap, see e.g. Ref. [24], which has attracted significant attention across different research fields, including condensed matter (see Ref. [25] for a review) and quantum thermodynamics [26]. In this work, we provide a new exact solution for a driven quantum many-body system. Specifically, we solve the dynamics of a tight-binding chain under an arbitrary time-dependent linear potential, giving access to the dynamics of observables such as particle density, current, and of the entanglement entropy.

### v2

@@ 1 Introduction
In the equilibrium setting, progress has been made in related inhomogeneous models, such as the gradient XX chain studied in Refs. [6, 7, 8], which allows for an exact characterization of local observables and entanglement. In contrast, when the linear potential acquires a time-dependent slope, the physics becomes more involved. Some special cases are known—such as quenches in the gradient XX chain [8], traveling impurity [9, 10], or periodic driving [11, 12, 13, 14, 15, 16]. However, while hydrodynamic and numerical approaches have provided valuable insights [17, 18], a general exact solution for the dynamics of such driven Stark-localized systems has not been fully established. Time-dependent linear potentials can give rise to nontrivial effects that go beyond simple generalizations of the static case. Standard examples include the periodic modulation of the driving field which may lead to dynamic localization [19, 20], where the motion of a wavepacket is effectively frozen at certain parameter values of the periodic modulation, as well as related phenomena such as coherent destruction of tunneling [21, 3]. These examples illustrate that driving and localization may interplay in a nontrivial way and underscore the need to explore more general driving protocols.This goal is also motivated by the growing interest in quantum control [22, 23, 24], Floquet engineering [25, 26], and transport [27], where the resulting nonequilibrium dynamics can have rich and unexpected behavior.
More generally, and even restricting to non-interacting systems, exact results in the presence of driving are remarkably scarce [28]. Notable exceptions include the case of a hardcore Bose gas in a time-dependent harmonic trap, see e.g. Ref. [29], which has attracted significant attention across different research fields, including condensed matter (see Ref. [30] for a review) and quantum thermodynamics [31]. In this work, we provide a new exact solution for a driven quantum many-body system. Specifically, we solve the dynamics of a tight-binding chain under an arbitrary time-dependent linear potential, giving access to the dynamics of observables such as particle density, current, and of the entanglement entropy.